OpenStudy (anonymous):

Verify the identity. \[\cot \left( x-\frac{ \pi }{ 2 } \right)=-tanx\]

OpenStudy (anonymous):

I don't know if I should use cofunction identites because cot (pi/2-x) = tan x, but there's a negative sign.

OpenStudy (anonymous):

\[\cot \left( x-\frac{ \pi }{ 2 } \right)=-tanx\]

OpenStudy (lynfran):

use compound identity

OpenStudy (anonymous):

What is that?

OpenStudy (lynfran):

\[\frac{ \cos(x-\frac{ \pi }{ 2 }) }{ \sin(x-\frac{ \pi }{ 2 }) }\]

OpenStudy (anonymous):

Oh okay I get it now. Thank you!

OpenStudy (lynfran):

\[\frac{ cosxcos \frac{ \pi }{ 2 }+sinxsin \frac{ \pi }{ 2 } }{ sinxcos \frac{ \pi }{ 2 }-cosxsin \frac{ \pi }{ 2 } }\]

OpenStudy (lynfran):

\[\frac{ cosx(0)+sinx(1) }{sinx(0)-cosx(1) }\]

OpenStudy (welshfella):

= - tan x yes good work LynFran